OpenStudy (anonymous):

evaluate. 64^2/3 / 216 ^-1/3

OpenStudy (anonymous):

hint 4^3=64 and 6^3=216

OpenStudy (anonymous):

okay so 4^2 / 6^1 ? 16/6 ? :S

OpenStudy (anonymous):

close but you 16/(1/6) = 16*6=?

OpenStudy (anonymous):

96. thank you ! :)

OpenStudy (anonymous):

:-)